Suppose a recurrence relation


an=7an−1−12an−2

where a1=16 and a2=52


can be represented in explicit formula, either as:


Formula 1:

an=pxn+qnxn

                      or  


Formula 2:

an=pxn+qyn

 


where 

x

and

y

are roots of the characteristic equation.


**If the explicit formula is in the form of Formula 2, consider p < q.


Determine

p and q




1
Expert's answer
2021-12-02T06:57:39-0500

 characteristic equation:

"x^2-7x+12=0"

"x=\frac{7\pm\sqrt{49-48}}{2}"

"x_1=3,x_2=4"


solution:

"a_n=p\cdot 3^n+q\cdot 4^n"

"a_1=3p+4q=16"

"a_2=9p+16q=52"


"4q=4 \implies q=1"

"p=(16-4q)\/3=4"